# Robust purely optical signatures of Floquet-Bloch states in crystals

### v1

@@ Abstract
Strong light-matter interactions enable triggering on-demand novel functionalities in materials. For periodically driven solids, the Floquet-Bloch states have emerged as the natural states to characterize the physical properties of these non-equilibrium systems. However, robust signatures of the Floquet-Bloch states are needed as common experimental conditions, such as pulsed laser excitation and dissipative many-body dynamics can disrupt their formation and survival. Here, we identify a tell-tale signature of Floquet-Bloch states in the linear optical response of laser-dressed solids that remains prominent even in the presence of strong spectral congestion of bulk matter. To do so, we introduce a computationally efficient strategy to finally capture the response properties of realistic laser-dressed solids, and use it to compute the optical absorption of ZnO using a model of full-dimensionality. The computations reveal intense, spectrally isolated, absorption/stimulated emission features at mid-infrared energies present for a wide range of laser-driving conditions that arise due to the hybridization of the Floquet-Bloch states. As such, these spectral features open a purely optical pathway to investigate the birth and survival of Floquet-Bloch states while avoiding the experimental challenges of fully reconstructing the band structure.
@@
Because experimental realizations of Floquet engineering rarely satisfy Floquet theorem –that requires unitary dynamics under strict periodic driving– there has been a significant recent interest in identifying tell-tale signatures of the Floquet-Bloch states that can signal their birth and evolution under effects such as pulsed lasers and dissipative many-body dynamics that can disrupt their formation and survival. For this, time- and angle-resolved photoemission spectroscopy (Tr-ARPES) has emerged as a useful technique [14] to reconstruct the band structure and visualize the characteristic Floquet replicas [9, 15, 16, 17, 18, 19].
@@
As an alternative, purely optical absorption experiments would be preferred as they are more common and often easier to realize experimentally. In the optical response, laser-dressing leads to absorption sidebands due to the Floquet replicas that have been computationally illustrated in model one-dimensional solids [13] and experimentally demonstrated in systems of low dimensionality [20, 21, 22, 23, 24, 25]. However, in bulk materials, although the Floquet replicas have been observed in Tr-ARPES [17, 16, 18, 19], they have remained elusive in the optical response [26, 27, 10, 28, 29], presumably because they are obscured by spectral congestion and strong exciton transitions. This has prevented the use of purely optical signatures to study the emergence and evolution of Floquet states in materials.
Systematic progress requires understanding the response properties of laser-dressed solids using atomistically detailed models. However, this has remained a formidable challenge because of the computational cost of propagating the quantum dynamics given the large (s) number of bands needed for convergence [30, 31, 32] and the challenges of properly sampling reciprocal space and the frequency dependence of the response in three-dimensional solids [33, 34, 35, 36, 37]. For this reason, theoretical simulations often resort to simple few-band models of reduced dimensionality [38, 39, 40, 41] that are not necessarily representative of experimentally relevant systems.
@@
In this Letter, we demonstrate the first computations of the optical absorption spectra of a realistic laser-dressed three-dimensional solid –ZnO described using a density functional theory (DFT) Hamiltonian. As detailed in our companion paper [42], these computations are now feasible due to the development of: (i) a response theory of laser-dressed solids; (ii) a truncated velocity gauge that has highly favorable convergence properties with the number of bands; (iii) the massively parallel computational implementation of the theory and; (iv) its integration with electronic structure codes for solids [43]. Our computations reveal robust, spectrally isolated, intense optical features at mid-infrared frequencies that provide a purely optical tell-tale signature of the emergence of Floquet states. These features can be used to monitor their birth and evolution in solids.
We simulate a physical situation in which a continuous wave laser of arbitrary intensity and frequency drives a solid of arbitrary band-structure and dimensionality out of equilibrium. The effective absorption properties of this laser-dressed solid are then probed by a weak continuous wave laser source. To capture the light-matter interaction, we introduce the truncated velocity gauge [44, 42], where is expressed as a series of nested commutators to a given order. Here, is the position and the Hamiltonian operator of the -electron system, its conjugate momentum operator, the vector potential of light, the electron’s charge and its mass. As detailed in Ref. [42], like the usual velocity gauge, this structure has the advantage of respecting the space-periodicity in solids as needed to invoke Bloch theorem; yet it overcomes its poor convergence properties [30, 31, 32] with the number of bands and its incompatibility with DFT Hamiltonians that employ non-local pseudopotentials. For computations, we consider generalized tight-binding models constructed from Wannier interpolation of the DFT electronic structure as it allows computing the nested commutators to any order. Overall, the truncated velocity gauge enables tractable converged computations of Floquet engineering in realistic solids.
@@
To capture the optical absorption, we quantify the rate at which the probe laser leads to transitions among the laser-dressed states [45, 13]. The total Hamiltonian is , where describes the material and its interaction with a drive laser with period , amplitude and polarization . In turn, describes the interaction with the probe laser of amplitude , frequency and polarization . We consider the effect of the drive laser exactly while the influence of the probe is captured to first-order in perturbation theory. In second quantization, , where is the single-particle laser-dressed Hamiltonian. Here, creates (annihilates) a fermion in Bloch state with band index , and crystal momentum in the first Brillouin zone (BZ), where is the Bloch mode and the crystal’s volume. Since is time-periodic, we can invoke Floquet theorem where now the Floquet-Bloch states are solutions to the time-dependent Schrödinger equation . The quasienergies and Floquet-Bloch modes are solutions to the eigenvalue problem in Sambe space [46]. The quasienergies are uniquely defined in the first Floquet-Brillouin zone (FBZ) such that for BZ. The Floquet-Bloch modes are time-periodic with period and space-periodic with period –the lattice vector of the crystal.
@@
The optical absorption coefficient of a laser-dressed solid at probe photon energy is [42]:
@@
Equation (1) is valid for any arbitrary time-periodic driving laser and for any space-periodic material including semiconductors, semi-metals, insulators and quantum materials. Its structure is analogous to the near-equilibrium optical response [47] but with the Floquet-Bloch modes playing the role of the pristine eigenstates. Specifically, optical absorption in laser-dressed solid arises due to transitions among Floquet-Bloch modes separated by Floquet-Brillouin zones with Bohr transition energy . The first term represents absorption, the second stimulated emission. Transition intensities are determined by the population factor and the momentum matrix element . guarantees that an optical transition happens from an initially occupied state to an empty one, and depends on the initial band occupation and also on the drive. Equation 1 underscores that the Floquet-Bloch modes provide the natural basis to understand the response properties of laser-dressed solids.
To find optical signatures of Floquet-Bloch states, we use Eq. (1) to compute the absorption spectrum for laser-dressed wurtzite ZnO crystal using the computational strategy detailed in Ref. [42]. The DFT electronic structure of ZnO was computed using Quantum Espresso [48] with Hubbard (DFT+U) corrections needed to reproduce experimental band gaps [49, 50, 51] (technical details in the Supplementary Material (SM) [52]).
Since the computational implementation of Eq. 1 is based on maximally-localized Wannier functions, we first perform a Wannier interpolation of the electronic structure using Wannier90 [53]. Figure 1(a) compares the band structure along the high-symmetry points obtained using DFT+U and the Wannier Hamiltonian constructed using 22 Wannier functions. As shown, the 22-Wannier model accurately reproduces the DFT+U band structure in a wide energy window. Figure 1(b) shows the computed field-free optical absorption spectrum of ZnO with the BZ sampled using a dense -space grid. The spectrum shows a sharp band-edge at eV as observed experimentally [26, 51]. Increasing the number of Wannier functions leaves the band-edge unchanged. Throughout, spectral features are Lorentzian broadened with width 0.02 eV to account for the finite energy resolution of femtosecond Gaussian probe pulses [15]. Further, transitions with energy less than 0.06 eV are not reported to account for the limits of the finite BZ discretization. We take the drive and probe to be linearly polarized along the crystal’s -axis. Other polarization directions yield qualitatively similar physics.
@@
Figure 2 shows the optical absorption spectra of ZnO laser-dressed with field amplitudes V/Å (intensities 0.1-1.2 TW/cm2) in the intermediate regime of light-matter interaction (non-ionizing, but non-perturbative) and with frequency chosen to be near-resonance with the band-edge ( eV, Fig. 3(a)-(c)) and off-resonance ( eV, Fig. 3(d)-(f)). Computations employ 23 terms in the series of nested commutator needed for convergence.
@@
In the near-resonance case (fig. 3(a)-(c)), the laser-dressing strongly suppresses the absorption for and leaves the spectrum for eV unchanged. This suppression is due to band-gap renormalization created by hybridization of the Floquet-Bloch states (see associated Tr-ARPES evidence in Ref. [17]). More explicitly, when the solid is resonantly driven, the first Floquet-Bloch replica of the valence (conduction) band overlaps with the conduction (valence) band. This leads to band hybridization that opens energy gap at , thus removing transitions at this frequency. This band-gap renormalization is an optical signal of the formation of Floquet-Bloch states as it occurs due to their hybridization. However, the challenge with isolating this feature in the optical spectra is that it emerges for around the band-edge where it can be obscured by excitonic features [54, 55, 51] which are not included in the model.
@@
For off-resonance driving (Fig. 2(d)-(f)), the laser-dressing leads to below band-edge absorption features eV that spread to even lower energies as the field strength is increased. In a Floquet-Bloch picture, this so-called dynamical Franz-Keldysh Effect [38] can be attributed to transitions occurring among the Floquet-Bloch replicas of valence and conduction bands [13]. For off-resonance drive laser, multiple Floquet-Bloch replicas are formed leading to transitions with energies lower than the field-free band-edge. Further, since the replicas are separated from each other by the drive photon energy, they can lead to absorption sidebands. Both the below-band gap features and the sidebands are optical signatures of Floquet states. However, as the band renormalization, they are also challenging to observe. Below-band gap absorption overlaps in energy with dominant exciton features. In turn, absorption sidebands are obscured in Fig. 2(d)-(f) by the spectral congestion of the three-dimensional material.
@@
Interestingly, Fig. 2 show prominent absorption and stimulated emission features at mid-infrared frequencies ( eV) that emerge under both near-resonance (Fig. 2(a)-(c)) and off-resonance (Fig. 2(e)-(f)) driving. These features are well separated from the band-edge or vibrational features, and are thus not obscured by possible excitonic contribution or spectral congestion. Further, their emergence is robust to changes in the drive laser frequency and amplitude.
@@
To clarify their physical origin, we developed a minimal one-dimensional two-band model that captures this phenomenon from a projection of the highest valence and the lowest conduction band of ZnO constructed using perturbation theory [56] (see SM[52] for model parameters). This model system also shows low-frequency transitions for the same drive laser parameters as ZnO (see SM[52] for the absorption spectra computed using the method in Ref. [13]). However, its reduced dimensionality enables to more clearly identify the essential physics.
@@
As we now show, these low-frequency features arise due to formation and subsequent hybridization of the Floquet-Bloch modes. Figure 3 top row shows this for drive laser frequency near-resonant to the band-edge. We plot the quasienergies of the two-band model within a FBZ for a representative -point ( units of , with Å the unit cell length) in Fig. 3(a) as a function of . The two quasienergies form an avoided crossing around eV which is a signature of level hybridization. Absorption spectrum of this two-band model with driving frequency chosen before the avoided crossing ( eV purple vertical line in Fig. 3(a)), shown in Fig. 3(c) shows no significant low-frequency feature. But when the model is driven by laser parameters chosen at the avoided crossing ( eV green vertical line in Fig. 3(a)), the spectrum (Fig. 3(e)) exhibits intense low-frequency absorption feature centered at eV with strong contribution from the transition at (blue line).
@@
Similar behavior is also seen for off-resonance drive laser with eV. The two quasienergies in a FBZ for the same representative as a function of in Fig. 3(b) show an avoided crossing centered around V/Å. When the model is driven by laser with V/Å before the avoided crossing (purple vertical line in Fig. 3(b)), the spectrum shown in Fig. 3(d) shows no significant low-frequency features in general and faint transition for in particular (blue line). But when the system is driven by a laser with V/Å at the avoided crossing (green vertical line in Fig. 3(b)), the spectrum shown in Fig. 3(f) shows intense absorption and stimulated emission features. This shows that in both resonantly and non-resonantly driven solid, the hybridization of the Floquet-Bloch modes indicated by the opening of avoided crossings induces low-frequency transition signals.
@@
Figures 3(d) and (h) summarize the emergent physical picture. Resonant and off-resonance laser-dressing leads to formation of Floquet-Bloch replicas (blue lines) of the valence and conduction band (black lines) that can energetically overlap either with the original Bloch bands or their replicas. This induces hybridization of the Floquet-Bloch states and formation of hybrid bands (red lines) now with avoided crossings as a function of and . Probing this laser-dressed band-structure leads to intense low-frequency transitions as indicated by purple arrows in Fig. 3(d) for resonant and Fig. 3(h) for off-resonance case.
@@
This origin of the low-frequency features in the two-band model can be naturally extended to laser-dressed ZnO. Figure 3 shows that for a single -point, the two Floquet-Bloch modes form an avoided crossing in the band structure due to hybridization for certain drive laser parameters. In the realistic computations, similar hybridization occurs but for a large number of -vectors given that a sufficient driving laser amplitude is applied. This leads to a plethora of low-frequency transitions, some that lead to absorption while others lead to stimulated emission depending on whether the laser parameters are before or after each specific avoided crossing. These contributions do not exactly cancel one another leading to the robust low-frequency features observed in Fig. 3.
These low-frequency features are a clear indication of the formation Floquet-Bloch states in laser-dressed materials. Our computations suggest that to observe them in semiconductors requires moderate field intensities ( TW/cm2) for lasers with frequency near-resonant to the band-edge or stronger lasers ( TW/cm2) if dressing off-resonance. We expect these signals to survive scattering effects as the Floquet-Bloch states form following the completion of just one laser pulse cycle in few-cycle lasers [15]. For the laser frequencies considered here, this corresponds to a maximum time scale of 10 fs, which is significantly shorter than the typical 20-200 fs semiconductor [57] scattering time. Few-cycle Gaussian probe pulses with energy resolution limit of eV (corresponding to time resolution limit of 17 fs) needed to conveniently detect the low-frequency signals have been recently employed in experiment [15]. Thus, the observation of these features is within the reach of present-day experiments.
@@
To summarize, we presented the first computational demonstration of the optical absorption spectrum of a laser-dressed solid described using realistic DFT electronic structure of full dimensionality. We found that the absorption spectra of a laser-dressed ZnO crystal predicts prominent absorption/stimulated emissions at mid-infrared frequencies that are robust to drive laser parameters, survive the spectral congestion in solids, and are well isolated from the field free band-edge transitions and possible excitonic features. Using a minimalistic two-band model, we showed that these features arise due to the hybridization of the Floquet-Bloch modes, time-and space-periodic functions that define the Floquet-Bloch states. We propose that these features provide a robust purely optical tell-tale signature of the formation of Floquet-Bloch states in bulk matter needed to study Floquet engineering in realistic conditions through simple electroabsorption spectroscopy, thus avoiding full band-structure reconstruction.

### v2

# Robust purely optical signatures of Floquet states in laser-dressed crystals

@@ Abstract
Strong light-matter interactions can create non-equilibrium materials with on-demand novel functionalities. For periodically driven solids, the Floquet theorem provides the natural states to characterize the physical properties of these laser-dressed systems. However, signatures of the Floquet states are needed, as common experimental conditions, such as pulsed laser excitation and dissipative many-body dynamics, can disrupt their formation and survival. Here, we identify a tell-tale signature of Floquet states in the linear optical response of laser-dressed solids that remains prominent even in the presence of strong spectral congestion of bulk matter. To do so, we introduce a computationally efficient strategy based on the Floquet formalism to finally capture the full frequency-dependence in the optical response properties of realistic laser-dressed crystals, and use it investigate the Floquet engineering in a first-principle model for ZnO of full dimensionality. The computations reveal intense, spectrally isolated, laser-controllable, absorption/stimulated emission features at mid-infrared energies present for a wide range of laser-driving conditions that arise due to the hybridization of the Floquet states. As such, these spectral features open a purely optical pathway to investigate the birth and survival of Floquet states while avoiding the experimental challenges of fully reconstructing the band structure.
@@
Because experimental realizations of Floquet engineering rarely satisfy Floquet theorem –that requires unitary dynamics under strict periodic driving– there has been a significant recent interest in identifying tell-tale signatures of the Floquet-Bloch states that can signal their birth and evolution under effects such as pulsed lasers and dissipative many-body dynamics that can disrupt their formation and survival. For this, time- and angle-resolved photoemission spectroscopy (Tr-ARPES) has emerged as a useful technique [14] to reconstruct the band structure and visualize the characteristic Floquet replicas [9, 12, 15, 16, 17, 18, 19], distorted copies of the electronic states separated in energy by multiples of the driving laser photon energy.
@@
As an alternative, purely optical absorption experiments would be preferred as they are more common and often easier to realize experimentally. In the linear optical response of laser-dressed matter, the Floquet replicas lead to absorption sidebands that have been computationally illustrated in model one-dimensional solids [13] and experimentally demonstrated in systems of low dimensionality[20, 21, 22, 23, 24, 25]. However, in bulk materials, although the Floquet replicas have been observed in Tr-ARPES [17, 16], they have remained elusive in the optical response, presumably because they are obscured by spectral congestion. This has prevented the use of optical absorption spectroscopy to study the emergence and evolution of Floquet states in materials.
Systematic progress requires developing a non-equilibrium theory of the linear response properties of laser-dressed solids and quantifying such a response in atomistically detailed models. According to the fluctuation-dissipation theorem, for near-equilibrium matter the response at frequency is dictated by the Fourier transform of one-time correlation functions of the thermal state [26, 27, 28]. By contrast, for non-equilibrium laser-dressed materials, the response at frequency is now determined by a double Fourier transform of two-time correlation functions because the time-translational symmetry is broken. This quantity is challenging to evaluate as it requires propagating the many-body quantum dynamics for long times forward and back for each pair of times and and for each frequency [29, 13].
@@
In this Letter, we demonstrate the first computations of the non-equilibrium optical absorption spectra of a realistic laser-dressed three-dimensional solid –ZnO described using a first-principle density functional theory (DFT) Hamiltonian. As detailed in our companion paper [30], these computations are now feasible due to the development of: (i) a Floquet-based approach to efficiently compute the full frequency dependence of the optical response of laser-dressed solids; (ii) a truncated velocity gauge for the light-matter interactions that enables the use of first-principle Hamiltonians and overcomes well-known convergence issues of the velocity gauge with the number of bands [31, 32, 33]; (iii) the massively parallel computational implementation of the theory and; (iv) its integration with electronic structure codes for solids [34]. Our computations reveal robust, laser-controllable, spectrally isolated, intense optical features at mid-infrared frequencies that provide a purely optical tell-tale signature of the emergence of Floquet states. These features can be used to monitor their birth and evolution in solids through optical absorption spectroscopy.
We simulate a drive-probe physical situation in which a continuous wave laser of arbitrary intensity and frequency drives a solid of arbitrary band-structure and dimensionality out of equilibrium. The effective absorption properties of this laser-dressed solid are then probed by a weak continuous wave laser source. We treat the drive laser exactly through a Floquet formalism while the influence of the probe is captured to first-order in perturbation theory. In our companion paper [30], we detail the underlying theory and computational framework to compute the frequency-dependent optical absorption spectra of laser-dressed solids using first-principle Hamiltonians.
@@
Briefly, the total Hamiltonian is , where describes the material and its interaction with a drive laser with period , amplitude and polarization . In turn, describes the interaction with the probe laser of amplitude , frequency and polarization . We consider the effect of the drive laser exactly while the influence of the probe is captured to first-order in perturbation theory. In second quantization, , where is the single-particle laser-dressed Hamiltonian. Here, creates (annihilates) a fermion in Bloch state with band index , and crystal momentum in the first Brillouin zone (BZ), where is the Bloch mode and the crystal’s volume. Since is time-periodic, we can invoke Floquet theorem where now the Floquet-Bloch states are solutions to the time-dependent Schrödinger equation . The quasienergies and Floquet-Bloch modes are solutions to the eigenvalue problem in Sambe space [35]. The quasienergies are uniquely defined in the first Floquet-Brillouin zone (FBZ) such that for BZ. The Floquet-Bloch modes are time-periodic with period and space-periodic with period –the lattice vector of the crystal.
@@
The non-equilibrium optical absorption coefficient of a laser-dressed crystal at probe photon energy is [30]:
@@
Equation (1) is valid for any arbitrary time-periodic driving laser and for any space-periodic material including semiconductors, semi-metals, insulators and quantum materials. When interfaced with DFT computations, it captures electronic correlations at preparation time. Remarkably, its structure is analogous to the near-equilibrium optical response [36] but with the Floquet-Bloch modes playing the role of the pristine eigenstates. Specifically, optical absorption in laser-dressed solids arise due to transitions among Floquet-Bloch modes separated by Floquet-Brillouin zones with Bohr transition energy . Here where is the maximum number of FBZs needed for convergence. The first term represents absorption, the second stimulated emission. Transition intensities are determined by the population factor (that guarantees transitions occur from an occupied to an empty state) and the momentum matrix element (that are dependent on the number of FBZs separating the modes). Interestingly, in this non-equilibrium context, transition matrix elements and population factors depend on the strength and nature of the driving laser (instead of being properties inherent to the material only), and optical transitions occur among states that can be separated by a variable number of FBZs leading to absorption sidebands.
To find optical signatures of Floquet states, we use Eq. (1) to compute the absorption spectrum for laser-dressed wurtzite ZnO crystal using a first-principle Hamiltonian and the computational strategy detailed in Ref. [30]. The large eV ZnO band-gap enables resonant and non-resonant laser-dressing using experimentally accessible frequencies [37, 38]. The DFT electronic structure of ZnO was computed using Quantum Espresso [39] with Hubbard (DFT+U) corrections needed to reproduce experimental band gaps [40, 41, 42] followed by a Wannier interpolation using Wannier90 [43] that enables efficient computation of Eq. (1). The employed 22-Wannier model accurately reproduces the DFT band structure in a 25 eV energy range (see Fig. S1 in Supplemental Material (SM)[44]) and, together with and a dense -space grid grid of , provides converged results for the non-equilibrium optical absorption. Throughout, since electron-phonon interactions introduce Franck-Condon like progressions [45], in our computations spectral features are Lorentzian broadened with width 0.02 eV. Further, transitions with energy less than 0.06 eV are not reported to account for the limits of the finite BZ discretization. We take the drive and probe to be linearly polarized along the crystal’s -axis. Other polarization directions yield qualitatively similar physics. Additional computational details are included in the SM [44].
@@
Figure 1 shows the optical absorption spectra of pristine ZnO (gray lines) and laser-dressed (blue lines) ZnO with field amplitudes V/Å (intensities 0.1-1.2 TW/cm2) in the intermediate regime of light-matter interaction (non-ionizing, but non-perturbative) and with frequency chosen to be near-resonance with the band-edge ( eV, Fig.1a-c) and off-resonance ( eV, Fig.1d-f). The pristine spectrum shows a sharp band-edge at eV as observed experimentally [38, 42]. Floquet engineering leads to strong modification of the optical response.
@@
In the near-resonance case (Fig. 1a-c), the laser-dressing strongly suppresses the absorption for and leaves the spectrum for eV unchanged (feature I). This suppression is due to band-gap renormalization created by hybridization of the Floquet-Bloch states (see associated Tr-ARPES evidence in Ref. [17]). More explicitly, when the solid is resonantly driven, the first Floquet replica of the valence (conduction) band overlaps with the conduction (valence) band. This leads to band hybridization that opens energy gap at , thus removing transitions at this frequency. This band-gap renormalization is an optical signal of the formation of Floquet-Bloch states as it occurs due to their hybridization. However, the challenge with isolating this feature in the optical spectra is that it emerges for around the band-edge where it can be obscured by excitonic features [46, 47, 48] which are not included in the model.
@@
For off-resonance driving (Fig. 1d-f), the laser-dressing leads to below band-edge absorption features eV (II) that spread to even lower energies as the field strength is increased. This so-called dynamical Franz-Keldysh Effect [49] can be attributed to transitions occurring among the Floquet replicas of valence and conduction bands [13]. For off-resonance drive laser, multiple Floquet replicas are formed leading to transitions with energies lower than the field-free band-edge. Further, since the replicas are separated from each other by the drive photon energy, they can lead to absorption sidebands. Both the below-band gap features and the sidebands are optical signatures of Floquet states. However, as the band renormalization, they are also challenging to observe. Below-band gap absorption overlaps in energy with dominant exciton features. In turn, while the absorption sidebands are expected from Eq. (1), and clearly visible in systems of low-dimensionality [23] including one-dimensional model crystals [13], they are obscured in Fig. 1d-f by the spectral congestion of the three-dimensional material.
@@
Interestingly, Fig. 1 shows prominent absorption and stimulated emission features (III) at mid-infrared frequencies ( eV) that emerge under both near-resonance (Fig. 1a-c) and off-resonance (Fig. 1e-f) driving. These features are well separated from the band-edge or vibrational features, and are thus not obscured by possible excitonic contribution or spectral congestion. Further, their emergence is robust to changes in the drive laser frequency and amplitude.
@@
To clarify their physical origin, we developed a minimal one-dimensional two-band model that captures this phenomenon from a projection of the highest valence and the lowest conduction band of ZnO constructed using perturbation theory [50] (see SM[44] for model parameters). This model system also shows low-frequency transitions for the same drive laser parameters as ZnO (see Fig. S3 in the SM[44]). However, its reduced dimensionality enables to more clearly identify the essential physics.
@@
As we now show, these low-frequency features arise due to formation and subsequent hybridization of the Floquet-Bloch modes. Figure 2 top row shows this for drive laser frequency near-resonant to the band-edge. We plot the quasienergies of the two-band model within a FBZ for a representative -point ( units of , with Å the unit cell length) in Fig. 2a as a function of . The two quasienergies form an avoided crossing around eV which is a signature of level hybridization. Absorption spectrum of this two-band model with driving frequency chosen before the avoided crossing ( eV purple vertical line in Fig. 2a), shown in Fig. 2c (purple line) shows no significant low-frequency feature. But when the model is driven by laser parameters chosen at the avoided crossing ( eV green vertical line in Fig. 2a), the spectrum (Fig. 2c green line) exhibits an intense low-frequency absorption feature centered at eV with strong contribution from the transition at .
@@
Similar behavior is also seen for off-resonance drive laser with eV. The two quasienergies in a FBZ for the same representative as a function of in Fig. 2b show an avoided crossing centered around V/Å. When the model is driven by laser with V/Å before the avoided crossing (purple vertical line in Fig. 2b), the spectrum shown in Fig. 2d (purple line) shows no significant low-frequency features in general and faint transition for in particular. But when the system is driven by a laser with V/Å at the avoided crossing (green vertical line in Fig. 2b), the spectrum shown in Fig. 2d (green line) shows intense absorption and stimulated emission features. This shows that in both resonantly and non-resonantly driven solid, the hybridization of the Floquet-Bloch modes indicated by the opening of avoided crossings induces low-frequency transition signals.
@@
Figure 3a-b summarize the emergent physical picture. Laser-dressing leads to the formation of Floquet replicas (blue lines) of the valence and conduction band (black lines) that can energetically overlap either with the original Bloch bands or their replicas. This induces hybridization of the Floquet-Bloch states and formation of hybrid bands (red lines) that display avoided crossings as a function of laser parameters (Fig. 2a-b). Probing this laser-dressed band-structure leads to intense low-frequency transitions as indicated by the purple arrows. For the resonant case Fig. 3a, the low-frequency features arise because of intra FBZ ( in Eq. (1)) optical transitions. In turn, for the non-resonant case Fig. 3b, such features can arise because of intra () and inter FBZ () optical transitions. For the inter FBZ transition leads to absorption and leads to stimulated emission resulting in the characteristic double peak in Fig. 2d around (see also Fig. 1e), with the lower frequency peak yielding absorption and the higher-frequency one stimulated emission. Higher order transitions are possible but less prominent in the spectra for this laser-dressing. All these features emerge because of the hybridization of Floquet-Bloch modes, see Fig. 2.
Figure 3a-b suggests that these low-frequency features are controllable simply by varying the parameters of the driving laser. This is, in fact, clearly shown in computations in the two-band model shown in Fig. 3c-d. For resonant driving (Fig. 3c), increasing the laser amplitude blue-shifts the low-frequency features in the spectra (blue arrow). This is because the gap due to Floquet hybridization increases with the strength of the light-matter interactions, akin to the Autler-Townes effect [51, 52]. In turn, for non-resonant driving (Fig. 3d), the low-frequency feature due to inter FBZ transition () blue-shift upon increasing the drive laser frequency. This controllability can be used to characterize the emergent low-frequency features, and enable their resolution with ultrafast probes.
@@
This origin of the low-frequency features in the two-band model can be naturally extended to the realistic laser-dressed ZnO. Figure 2 shows that for a single -point, the two Floquet-Bloch modes form an avoided crossing in the band structure due to hybridization for certain drive laser parameters. In the realistic computations, similar hybridization occurs but for a large number of -vectors given that a sufficient driving laser amplitude is applied. This leads to a plethora of low-frequency transitions, some that lead to absorption while others lead to stimulated emission depending on whether the laser parameters are before or after each specific avoided crossing. These contributions do not exactly cancel one another leading to the robust low-frequency features observed in Fig. 2.
In nature, these low frequency features will arise from the hybridization of many-body Floquet replicas. For this reason, as explicitly demonstrated for a Hubbard-Holstein model [53, 54, 55, 56, 54] in Figs. S4-S7 in the SM [44] using exact diagonalization techniques [57, 58, 59, 60], electron correlations and electron-phonon couplings influence the frequency and magnitude of these transitions as they modulate the many-body Floquet states but do not obscure the effect. Further, the effect is robust to spectral broadening. In fact, as shown in Fig. S8 in the SM[44], increasing spectral broadening in a wide range of energies (0.02-0.2 eV) reduces but does not obscure the main low-frequency features.
These low-frequency features are a clear indication of the formation of Floquet states in laser-dressed materials. Our computations suggest that to observe them in semiconductors requires moderate field intensities ( TW/cm2) for lasers with frequency near-resonant to the band-edge or stronger lasers ( TW/cm2) if dressing off-resonance.
Dissipation, pulsed excitation and other effects that limit the applicability of the Floquet theory are expected to reduce the visibility of these low frequency features. Thus, the onset and disappearance of these laser-controllable low-frequency features can be used to test the applicability of Floquet engineering in crystals under experimentally relevant conditions. In particular, we expect that the optical controllability of these low-frequency features can be employed to characterize timescales of effects that break the Hamiltonian’s time-periodicity.
@@
To summarize, we presented the first computational demonstration of the non-equilibrium optical absorption spectrum of a laser-dressed solid (ZnO) described using first-principle electronic structure of full dimensionality. We show that the Floquet engineering can dramatically change the optical response of a crystal. In particular, we find prominent absorption/stimulated emission features at mid-infrared frequencies that are robust to drive laser parameters, survive the spectral congestion and broadening in solids, and are well isolated from the field-free band-edge and possible excitonic transitions. Using a minimalistic two-band model, we demonstrated that these features arise due to the hybridization of the Floquet states. We propose that these features provide a robust, laser-controllable, purely optical tell-tale signature of the formation of Floquet states in non-equilibrium bulk matter that can be used to study the validity of Floquet engineering with varying experimental conditions (i.e. temperature, material, pulse width, etc.). As such, these spectral features open a purely optical pathway to investigate the birth and survival of Floquet states while avoiding the experimental challenges of fully reconstructing the band structure.